OpenStudy (anonymous):

Simplify: ((1/x+h)-(1/x))/h

OpenStudy (osanseviero):

[1/(x + h) - 1/x]/h. By multiplying the numerator and denominator by x(x + h): [1/(x + h) - 1/x]/h = {x(x + h) * [1/(x + h) - 1/x]}/[hx(x + h)] = [x - (x + h)]/[hx(x + h)] = (x - x - h)/[hx(x + h)] = -h/[hx(x + h)] = -1/[x(x + h)]. http://answers.yahoo.com/question/index?qid=20110122181251AAAr2zz

OpenStudy (anonymous):

Again do the top part which results "h" since 1/x cancels each other then h divided by h is 1 :D